OpenStudy (anonymous):

where is y=sqrt(x-1) differentiable?

OpenStudy (anonymous):

Do you know what differentiable means?

OpenStudy (anonymous):

yes, where the function is continuous

geerky42 (geerky42):

And where is function continuous?

OpenStudy (anonymous):

x is greater than or equal to 1?

jimthompson5910 (jim_thompson5910):

y = sqrt(x-1) y = (x-1)^(1/2) y' = (1/2)(x-1)^(-1/2) y' = 1/(2*(x-1)^(1/2)) y' = 1/(2*sqrt(x-1)) Notice how the domain of y' is the set of numbers that are greater than 1 So y = sqrt(x-1) is differentiable when x > 1

OpenStudy (anonymous):

so it's only greater than and not equal to 1?

jimthompson5910 (jim_thompson5910):

exactly, it's not differentiable at x = 1

jimthompson5910 (jim_thompson5910):

and if you graph y = sqrt(x-1), you'll see that the endpoint is at x = 1, which means there is no possible way to find the derivative here